Part B: Lets consider a town with a low-capacity town route (1500 vehicles per hour) and a high-capacity bypass (3000 v/h), as shown below. The bypass is longer but has higher geometric design standards. Lets assume that during morning peak, 2000 drivers approach the town and all of them want to use the shortest route, i.e. via town center. Not all drivers will be able to use this route, since it would become too congested even before its ultimate capacity. This means that many drivers would choose the bypass to avoid long queues and delays, until the network reaches equilibrium.

Bypass capacity 3000 v/h Town route capacity 1000 v/h) Under equilibrium conditions, traffic arranges itself in congested networks in such a way that no individual trip maker can reduce his/her travel costs by switching to another route. If all trip makers perceive costs in the same way (no stochastic effects), then, under equilibrium conditions, traffic arranges itself in congested networks such that all used routes between an O-D pair have equal and minimum costs while all unused routes have greater or equal costs -> Wardrops first principle. The flows on the two routes will satisfy Wardrops equilibrium when the corresponding costs are identical. The following stand Cb-15 + 0.005x Vb Ct 10+0.02 xV where: Cb, Vb Travel time (cost) and volume via the bypass G, vt Travel time (cost) and volume via the town route For the travel demand of 2000 vehicles, assign the traffic on alternative routes, so that to satisfy Wardrops first principle. Then try to formulate the equation estimating volumes in a generic form (without numbers for volume) What do you observe? Hint: You may equate the two equations; also, you know the total flows (V

PLEASE SHOW WORK. WILL UPVOTE WITHIN A DAY.
Answer #1

We know the equation those are :

1. Cb=15+0.005*Vb

2. Ct=10+0.02*Vt

For equilibrium we know that both the costs would be equal

So it becomes

15+0.005*Vb=10+0.02*Vt

Now we also know that

Vb +Vt = 2000

Vb = 2000-Vt

15+0.005(2000-Vt)=10+0.02*Vt

0.015Vt=25-10

Vt=1000 vehicle

Vb =2000-1000 = 1000 vehicle

We can see that half the traffic moves through town and half through bypass

So the general equation for traffic through each road is

Traffic through town = V/2

Traffic through bypass =V/2

Where V is the total traffic volume

Thanks do like the answer